OpenStudy (anonymous):

What is the solution of the system of equations? 9x+6y=0 4x+4y=-2 Solve by elimination.

OpenStudy (anonymous):

x=1 y=-3/2

OpenStudy (anonymous):

I need to know how to do it, not just the answer. How'd you get that?

OpenStudy (anonymous):

9x+6y=0 4x+4y=-2 18x+12y=0 12x+12y=-6 6x=6 x=1 4+4y=-2 4y=-6 y=-3/2

OpenStudy (anonymous):

I multiplied the top equation by -2 and the bottom equation by 3. So then I got -18x-12y=0 12x+12y=-6 So, the y's cancel and you get: -6x=-6 so, x=1 Then, you plug in x into one of the original equations to find y. I plugged it into the 2nd equation since it has smaller numbers. 4(1)+4y=-2 ----> 4+4y=-2. Subtract the 4 on both sides. ----> 4y=-6 Divide 4 on both sides and you get y=-6/4 which reduces to -3/2 So the solution is (1, -3/2)